# SimWorld-Robotics: Synthesizing Photorealistic and Dynamic Urban Environments for Multimodal Robot Navigation and Collaboration

### v1

@@ 1 Introduction
There has been tremendous progress in engineering general-purpose robotics that can follow human instructions and perform open-ended tasks [anderson2018visionandlanguagenavigationinterpretingvisuallygrounded, NEURIPS2024_d8087dea, gan2021threedworldtransportchallengevisually, pmlr-v205-li23a, Shridhar_2020_CVPR], thanks to the advances in robot foundation models. Training these models requires a large amount of data, much of which can be generated in high-fidelity embodied simulators, such as Habitat 3 [puig2023habitat30cohabitathumans], RoboTHOR [deitke2020robothoropensimulationtorealembodied], TDW [gan2021threedworldplatforminteractivemultimodal], VirtualHome [puig2018virtualhomesimulatinghouseholdactivities], Virtual Community [zhou2025virtualcommunityopenworld] and BEHAVIOR [pmlr-v205-li23a]. They can also be systematically evaluated in diverse scenarios created in these simulators. However, current embodied simulators for robotics have been focused on tabletop [mu2021maniskillgeneralizablemanipulationskill, yin2025partinstruct, mees2022calvinbenchmarklanguageconditionedpolicy, jiang2023vimageneralrobotmanipulation, zhang2023lohoravenslonghorizonlanguageconditionedbenchmark] or household tasks [szot2022habitat20traininghome, pmlr-v205-li23a, li2021igibson20objectcentricsimulation, srivastava2021behaviorbenchmarkeverydayhousehold, Shridhar_2020_CVPR]. In this work, we want to study how to create a realistic and scalable embodied simulator for outdoor robotics tasks.
@@
There have been urban simulators developed in recent years. However, to address the critical challenges faced by real-world robotics in urban environments, they lack the necessary realism, customizability, scalability, and versatility. For instance, well-known simulators such as AirSim [shah2017airsimhighfidelityvisualphysical], CARLA [Dosovitskiy17] mainly focus on autonomous driving domains. While it supports some manual customization to the provided city environments, it does not support procedural city environment generation. It also does not support the flexible control of embodied agents (such as mobile robots or pedestrians) other than vehicles. More recent city simulators, such as MetaDrive [li2021metadrive], MetaUrban [wu2024metaurbanembodiedaisimulation], significantly improve the scalability. However, the simulated environments still lack photorealism as shown in Figure 2.
@@ 2 Related Work
Embodied Simulators for Urban Environments. Recent embodied simulators mostly focus on the indoor environment, mainly designed for household robots [chang2017matterport3dlearningrgbddata, tsoi2022sean2, puig2018virtualhomesimulatinghouseholdactivities, puig2020watch, gan2021threedworldplatforminteractivemultimodal, puig2023habitat30cohabitathumans, kolve2022ai2thorinteractive3denvironment, pmlr-v205-li23a]. There have been outdoor simulators [Dosovitskiy17, shah2017airsimhighfidelityvisualphysical, wu2024metaurban, gao2024embodiedcitybenchmarkplatformembodied, wang2024grutopiadreamgeneralrobots, fan2023aerialvisionanddialognavigation] that provide more open-ended tasks, facilitate more complicated dynamics, but face challenges in reality and multi-agent parallelism. CARLA[Dosovitskiy17] is a flexible simulator dedicated to autonomous driving. AirSim [shah2017airsimhighfidelityvisualphysical] focuses on simulation for drone control. EmbodiedCity [gao2024embodiedcitybenchmarkplatformembodied] provides an arena for embodied AI with realistic urban settings and enriched agent control. Grutopia [wang2024grutopiadreamgeneralrobots] offers a large-scale annotated scene dataset for general service robot tasks. AerialVLN [liu2023aerialvlnvisionandlanguagenavigationuavs] serves as a highly photorealistic urban landscape for UAV navigation. MetaUrban [wu2024metaurban] features a wide range of micromobility tasks on the pedestrian, creating a highly complex urban environment. However, the dynamic scene interaction of human agents and social navigation (e.g., traffic rule following and pedestrian avoidance) has been much less explored.
@@
Instruction Following Benchmarks for Robot Navigation. Following complex instructions in the real world often requires understanding both language and vision cues in tandem. As summarized in Table 8 in Appendix B.1, prior benchmarks on vision-and-language navigation have tackled instruction following, but typically in static indoor scenes or street panoramas using only textual descriptions[pérezdarpino2020robotnavigationconstrainedpedestrian]. For instance, Touchdown [chen2020touchdownnaturallanguagenavigation] explores city navigation via language alone in a static street view setting, and R2R [anderson2018vision] focuses on multi-step instructions in household tasks. These settings lack moving obstacles and additional visual guidance, making them only partial proxies for real-world urban navigation. In contrast, our multimodal robot navigation benchmark, SimWorld-MMNav, challenges a robot to interpret and follow multimodal instructions (paired language instructions and visual hints) to reach the target location in a large-scale, photorealistic, dynamic city environment.
Multi-Robot Collaboration Benchmarks. There have been many recent multi-robot collaboration benchmarks as summarized in Table 9, but they do not evaluate multi-robot collaboration and communication for exploration and navigation in large urban environments. RoCo [beck2023roco] studies cooperative manipulation on tabletop tasks, without the need for environmental exploration. DOROTHIE [ma2022dorothie] introduces spoken dialogue for an ego vehicle, yet involves just one controllable agent. DriVLMe [driVLMe2023] simulates city-scale traffic with many vehicles, but lacks any verbal communication between robots. RobotSlang [kapoor2024robotslang] evaluates robot communication but is restricted to small environments. Where Are You? [das2018where] frames localization as a two-party dialogue, but the robots do not physically navigate to meet each other through a large environment.
@@
CARLA [Dosovitskiy17] $\times$ $\checkmark$ $\times$ 2 $\times$ $\checkmark$ $\checkmark$ $\checkmark$
AirSim [shah2017airsimhighfidelityvisualphysical] $\times$ $\checkmark$ $\times$ 2 $\times$ $\checkmark$ $\checkmark$ $\checkmark$
MetaUrban [wu2024metaurbanembodiedaisimulation] $\checkmark$ $\times$ $\times$ 2 $\checkmark$ $\times$ $\times$ $\checkmark$
EmbodiedCity [gao2024embodiedcitybenchmarkplatformembodied] $\times$ $\times$ $\times$ 2 $\checkmark$ $\times$ $\times$ $\times$
Grutopia [wang2024grutopiadreamgeneralrobots] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
AerialVLN [liu2023aerialvlnvisionandlanguagenavigationuavs] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
UnrealZoo [zhong2024unrealzoo] $\times$ $\checkmark$ $\checkmark$ 10 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ 3 SimWorld-Robotics
Built upon SimWorld [simworld], SimWorld-Robotics (SWR) introduces key extensions including procedural city generation, a traffic system, and support for an additional embodied agent: the quadruped robot. We begin by describing how SWR procedurally generates diverse and scalable urban environments with varying specifications. We then introduce the embodied agents supported in these environments—vehicles, humans, and robots—and explain the logic behind asynchronously controlling multiple agents. Finally, we outline the rule-based system governing background pedestrians and traffic dynamics. Details of SWR can be found in Appendix A.
@@ 3.2 Embodied Agents
SWR supports three types of embodied agents—humans, vehicles, and robots. Unlike synchronous designs [puig2018virtualhomesimulatinghouseholdactivities] where agents must wait for others to complete their actions, SWR allows asynchronous control, enabling each agent to act independently. To better reflect real-world conditions and support diverse tasks, SWR also provides rich action spaces for different agents and flexible observation spaces. The creation of new embodied agents in Appendix A.8.
Types of Embodied Agents. Unlike previous simulators [Dosovitskiy17, wu2024metaurban, li2021metadrive]that focus on a specific type of embodied agent—such as autonomous vehicles, robots, or humans—and mainly support control over that particular agent type, our simulator simultaneously supports all three major categories. This unified design enables the development of broader and more diverse embodied AI tasks within a single environment. In SWR, we have included all three types of embodied agents. For the robots, we have two kinds: the scooter and the quadruped robot.
@@ 3.3 Pedestrian and Traffic Simulation
The traffic simulation in SWR creates dynamic urban scenarios by orchestrating vehicle and pedestrian movement across a generated city map. It supports route assignment, intersection control, and pedestrian flow simulation, running on a fixed-time update loop for consistent real-time updates. Vehicle motion is governed by a feedback-based model using a PID controller, with empirically tuned parameters for realistic acceleration, braking, and turning dynamics [jain2024comprehensive]. Pedestrians follow a lightweight model, adjusting orientation incrementally toward their goals based on angular differences. To simulate realistic patterns, SWR uses a probabilistic routing strategy at intersections, where agents select paths based on predefined probabilities. This stochastic behavior introduces natural variability and enhances scene diversity. Details of predestrain and traffic simulation can be found in Appendix A.7.
@@ 4.1 Setup
SimWorld-20k. Based on SWR, we construct the SimWorld-20k dataset, including 100 training maps with an average area of $2\,\mathrm{km}^{2}$ and 200 oracle trajectories generated by A* [Hart1968] for 200 training tasks, whose average length is greater than $2.5\,\mathrm{km}$. Each trajectory contains over 100 steps, forming a training dataset of 20K steps. Details of it can be found in Appendix C.4.
@@
Evaluation Metrics. We evaluate navigation performance in the easy setting through three key metrics: (1) Success Rate (SR) [anderson2018evaluationembodiednavigationagents] measures the percentage of goal arrivals, (2) Subtask SR [Shridhar_2020_CVPR] tracks the ratio of completed subtasks, (3) Distance Progress [thomason2019visionanddialognavigation] evaluates instruction-following by the relative reduction in distance to the goal. The hard evaluation setting, which introduces a comprehensive traffic system and pedestrians, requires three additional safety metrics: (1) Static Collision counts collision between the robot and immovable objects like buildings and trees, (2) Dynamic Collision counts collisions with moving elements, such as pedestrians and vehicles, (3) Traffic Light Violation records the number of times the robot fails to adhere to traffic signal regulations.
@@
Baselines. Inspired by the recent success of large vision-language models (VLMs) on navigation tasks [wang2024qwen2], we evaluate multiple recent VLMs as backbones using ReAct [yao2023reactsynergizingreasoningacting], including GPT-4o, GPT-o3, GPT-o3-pro [4o], Gemini 2.5 Flash [gemini], Qwen-VL 2.5 [Qwen2.5-VL], Gemma 3 [gemma_2025] and InternVL [internvl2.5]. Additionally, we finetune QwenVL2.5-7B on SimWorld-20k. We also test a hybrid baseline, HybridGPT, where GPT-4o is used as a high-level decision maker and A* [Hart1968] is used as a low-level motion planner. For RL baselines, we train a multimodal policy model, VLA-RL, with DeBERTa-v3 [he2023debertav3improvingdebertausing] as language encoding and DINOv2 [oquab2024dinov2learningrobustvisual] as visual encoding, following VLN-CE [krantz2020navgraphvisionandlanguagenavigationcontinuous]. We include more baseline implementation details in Appendix C.5.
@@ 5.1 Setup
Evaluation Metrics. We assess multi-agent navigation performance through two principal metrics: (1) Collaborative Success Rate (CSR) [lowe2017multi] measures the percentage of tasks completed successfully before meeting; (2) Task Progress averages, across robots, the fraction of their shortest-path distance that is covered by the end of an episode.
Baselines. Following the collaboration paradigm of RoCo [beck2023roco], we enable robots to discuss a joint plan for concrete rendezvous behavior using a VLM. The follower robot will first describe its location using language for the main robot to localize it. Afterwards, two robots will confirm a plan, where the main robot will describe paths to a meeting location for the follower robot to follow, while the main robot will plan its own path to reach the meeting location. In SimWorld-MRS, we first evaluated GPT-4o and Gemini 2.5 Flash with oracal planner and then picked the best performing VLM (GPT-4o) and paired it with RoCo. We include more implementation details in the Appendix D.4.
@@ A.2 Full Comparison of SWR with Prior Simulators
Matterport3d [chang2017matterport3dlearningrgbddata] I $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Sean2.0 [tsoi2022sean2] I $\times$ $\times$ $\times$ 3 $\checkmark$ $\times$ $\times$ $\times$
@@
AI2THOR [kolve2022ai2thorinteractive3denvironment] I $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$
TDW [gan2021threedworldplatforminteractivemultimodal] I $\checkmark$ $\checkmark$ $\times$ 50 $\checkmark$ $\times$ $\times$ $\times$
SocNavBench [biswas2021socnavbenchgroundedsimulationtesting] I $\times$ $\times$ $\times$ 2 $\checkmark$ $\times$ $\times$ $\times$
Habitat 3.0 [puig2023habitat30cohabitathumans] I $\checkmark$ $\times$ $\checkmark$ 4 $\checkmark$ $\checkmark$ $\times$ $\checkmark$
VirtualHome 2.0 [puig2020watch] I $\checkmark$ $\times$ $\checkmark$ 25 $\times$ $\checkmark$ $\times$ $\times$
BEHAVIOR [pmlr-v205-li23a] I $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
CARLA [Dosovitskiy17] U $\times$ $\checkmark$ $\times$ 2 $\times$ $\checkmark$ $\checkmark$ $\checkmark$
AirSim [shah2017airsimhighfidelityvisualphysical] U $\times$ $\checkmark$ $\times$ 2 $\times$ $\checkmark$ $\checkmark$ $\checkmark$
MetaUrban [wu2024metaurbanembodiedaisimulation] U $\checkmark$ $\times$ $\times$ 2 $\checkmark$ $\times$ $\times$ $\checkmark$
EmbodiedCity [gao2024embodiedcitybenchmarkplatformembodied] U $\times$ $\times$ $\times$ 2 $\checkmark$ $\times$ $\times$ $\times$
Grutopia [wang2024grutopiadreamgeneralrobots] U $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
AerialVLN [liu2023aerialvlnvisionandlanguagenavigationuavs] U $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$

### v2

@@ 1 Introduction
There has been tremendous progress in engineering general-purpose robotics that can follow human instructions and perform open-ended tasks [3, 28, 17, 27, 44], thanks to the advances in robot foundation models. Training these models requires a large amount of data, much of which can be generated in high-fidelity embodied simulators, such as Habitat 3 [39], RoboTHOR [12], TDW [16], VirtualHome [37], Virtual Community [61] and BEHAVIOR [27]. They can also be systematically evaluated in diverse scenarios created in these simulators. However, current embodied simulators for robotics have been focused on tabletop [34, 57, 33, 23, 58] or household tasks [46, 27, 26, 45, 44]. In this work, we want to study how to create a realistic and scalable embodied simulator for outdoor robotics tasks.
@@
There have been urban simulators developed in recent years. However, to address the critical challenges faced by real-world robotics in urban environments, they lack the necessary realism, customizability, scalability, and versatility. For instance, well-known simulators such as AirSim [43], CARLA [14] mainly focus on autonomous driving domains. While it supports some manual customization to the provided city environments, it does not support procedural city environment generation. It also does not support the flexible control of embodied agents (such as mobile robots or pedestrians) other than vehicles. More recent city simulators, such as MetaDrive [29], MetaUrban [55], significantly improve the scalability. However, the simulated environments still lack photorealism as shown in Figure 2.
@@ 2 Related Work
Embodied Simulators for Urban Environments. Recent embodied simulators mostly focus on the indoor environment, mainly designed for household robots [8, 50, 37, 38, 16, 39, 24, 27]. There have been outdoor simulators [14, 43, 54, 18, 51, 15] that provide more open-ended tasks, facilitate more complicated dynamics, but face challenges in reality and multi-agent parallelism. CARLA[14] is a flexible simulator dedicated to autonomous driving. AirSim [43] focuses on simulation for drone control. EmbodiedCity [18] provides an arena for embodied AI with realistic urban settings and enriched agent control. Grutopia [51] offers a large-scale annotated scene dataset for general service robot tasks. AerialVLN [30] serves as a highly photorealistic urban landscape for UAV navigation. MetaUrban [54] features a wide range of micromobility tasks on the pedestrian, creating a highly complex urban environment. However, the dynamic scene interaction of human agents and social navigation (e.g., traffic rule following and pedestrian avoidance) has been much less explored.
@@
Instruction Following Benchmarks for Robot Navigation. Following complex instructions in the real world often requires understanding both language and vision cues in tandem. As summarized in Table 8 in Appendix B.1, prior benchmarks on vision-and-language navigation have tackled instruction following, but typically in static indoor scenes or street panoramas using only textual descriptions[pérezdarpino2020robotnavigationconstrainedpedestrian]. For instance, Touchdown [9] explores city navigation via language alone in a static street view setting, and R2R [2] focuses on multi-step instructions in household tasks. These settings lack moving obstacles and additional visual guidance, making them only partial proxies for real-world urban navigation. In contrast, our multimodal robot navigation benchmark, SimWorld-MMNav, challenges a robot to interpret and follow multimodal instructions (paired language instructions and visual hints) to reach the target location in a large-scale, photorealistic, dynamic city environment.
Multi-Robot Collaboration Benchmarks. There have been many recent multi-robot collaboration benchmarks as summarized in Table 9, but they do not evaluate multi-robot collaboration and communication for exploration and navigation in large urban environments. RoCo [59] studies cooperative manipulation on tabletop tasks, without the need for environmental exploration. DOROTHIE [32] introduces spoken dialogue for an ego vehicle, yet involves just one controllable agent. DriVLMe [21] simulates city-scale traffic with many vehicles, but lacks any verbal communication between robots. RobotSlang [5] evaluates robot communication but is restricted to small environments. Where Are You? [11] frames localization as a two-party dialogue, but the robots do not physically navigate to meet each other through a large environment.
@@
CARLA [14] $\times$ $\checkmark$ $\times$ 2 $\times$ $\checkmark$ $\checkmark$ $\checkmark$
AirSim [43] $\times$ $\checkmark$ $\times$ 2 $\times$ $\checkmark$ $\checkmark$ $\checkmark$
MetaUrban [55] $\checkmark$ $\times$ $\times$ 2 $\checkmark$ $\times$ $\times$ $\checkmark$
EmbodiedCity [18] $\times$ $\times$ $\times$ 2 $\checkmark$ $\times$ $\times$ $\times$
Grutopia [51] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
AerialVLN [30] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
UnrealZoo [60] $\times$ $\checkmark$ $\checkmark$ 10 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ 3 SimWorld-Robotics
Built upon SimWorld [42], SimWorld-Robotics (SWR) introduces key extensions including procedural city generation, a traffic system, and support for an additional embodied agent: the quadruped robot. We begin by describing how SWR procedurally generates diverse and scalable urban environments with varying specifications. We then introduce the embodied agents supported in these environments—vehicles, humans, and robots—and explain the logic behind asynchronously controlling multiple agents. Finally, we outline the rule-based system governing background pedestrians and traffic dynamics. Details of SWR can be found in Appendix A.
@@ 3.2 Embodied Agents
SWR supports three types of embodied agents—humans, vehicles, and robots. Unlike synchronous designs [37] where agents must wait for others to complete their actions, SWR allows asynchronous control, enabling each agent to act independently. To better reflect real-world conditions and support diverse tasks, SWR also provides rich action spaces for different agents and flexible observation spaces. The creation of new embodied agents in Appendix A.8.
Types of Embodied Agents. Unlike previous simulators [14, 54, 29]that focus on a specific type of embodied agent—such as autonomous vehicles, robots, or humans—and mainly support control over that particular agent type, our simulator simultaneously supports all three major categories. This unified design enables the development of broader and more diverse embodied AI tasks within a single environment. In SWR, we have included all three types of embodied agents. For the robots, we have two kinds: the scooter and the quadruped robot.
@@ 3.3 Pedestrian and Traffic Simulation
The traffic simulation in SWR creates dynamic urban scenarios by orchestrating vehicle and pedestrian movement across a generated city map. It supports route assignment, intersection control, and pedestrian flow simulation, running on a fixed-time update loop for consistent real-time updates. Vehicle motion is governed by a feedback-based model using a PID controller, with empirically tuned parameters for realistic acceleration, braking, and turning dynamics [22]. Pedestrians follow a lightweight model, adjusting orientation incrementally toward their goals based on angular differences. To simulate realistic patterns, SWR uses a probabilistic routing strategy at intersections, where agents select paths based on predefined probabilities. This stochastic behavior introduces natural variability and enhances scene diversity. Details of predestrain and traffic simulation can be found in Appendix A.7.
@@ 4.1 Setup
SimWorld-20k. Based on SWR, we construct the SimWorld-20k dataset, including 100 training maps with an average area of $2\,\mathrm{km}^{2}$ and 200 oracle trajectories generated by A* [19] for 200 training tasks, whose average length is greater than $2.5\,\mathrm{km}$. Each trajectory contains over 100 steps, forming a training dataset of 20K steps. Details of it can be found in Appendix C.4.
@@
Evaluation Metrics. We evaluate navigation performance in the easy setting through three key metrics: (1) Success Rate (SR) [1] measures the percentage of goal arrivals, (2) Subtask SR [44] tracks the ratio of completed subtasks, (3) Distance Progress [49] evaluates instruction-following by the relative reduction in distance to the goal. The hard evaluation setting, which introduces a comprehensive traffic system and pedestrians, requires three additional safety metrics: (1) Static Collision counts collision between the robot and immovable objects like buildings and trees, (2) Dynamic Collision counts collisions with moving elements, such as pedestrians and vehicles, (3) Traffic Light Violation records the number of times the robot fails to adhere to traffic signal regulations.
@@
Baselines. Inspired by the recent success of large vision-language models (VLMs) on navigation tasks [52], we evaluate multiple recent VLMs as backbones using ReAct [56], including GPT-4o, GPT-o3, GPT-o3-pro [35], Gemini 2.5 Flash [47], Qwen-VL 2.5 [4], Gemma 3 [48] and InternVL [10]. Additionally, we finetune QwenVL2.5-7B on SimWorld-20k. We also test a hybrid baseline, HybridGPT, where GPT-4o is used as a high-level decision maker and A* [19] is used as a low-level motion planner. For RL baselines, we train a multimodal policy model, VLA-RL, with DeBERTa-v3 [20] as language encoding and DINOv2 [36] as visual encoding, following VLN-CE [25]. We include more baseline implementation details in Appendix C.5.
@@ 5.1 Setup
Evaluation Metrics. We assess multi-agent navigation performance through two principal metrics: (1) Collaborative Success Rate (CSR) [31] measures the percentage of tasks completed successfully before meeting; (2) Task Progress averages, across robots, the fraction of their shortest-path distance that is covered by the end of an episode.
Baselines. Following the collaboration paradigm of RoCo [59], we enable robots to discuss a joint plan for concrete rendezvous behavior using a VLM. The follower robot will first describe its location using language for the main robot to localize it. Afterwards, two robots will confirm a plan, where the main robot will describe paths to a meeting location for the follower robot to follow, while the main robot will plan its own path to reach the meeting location. In SimWorld-MRS, we first evaluated GPT-4o and Gemini 2.5 Flash with oracal planner and then picked the best performing VLM (GPT-4o) and paired it with RoCo. We include more implementation details in the Appendix D.4.
@@ A.2 Full Comparison of SWR with Prior Simulators
Matterport3d [8] I $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Sean2.0 [50] I $\times$ $\times$ $\times$ 3 $\checkmark$ $\times$ $\times$ $\times$
@@
AI2THOR [24] I $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$
TDW [16] I $\checkmark$ $\checkmark$ $\times$ 50 $\checkmark$ $\times$ $\times$ $\times$
SocNavBench [6] I $\times$ $\times$ $\times$ 2 $\checkmark$ $\times$ $\times$ $\times$
Habitat 3.0 [39] I $\checkmark$ $\times$ $\checkmark$ 4 $\checkmark$ $\checkmark$ $\times$ $\checkmark$
VirtualHome 2.0 [38] I $\checkmark$ $\times$ $\checkmark$ 25 $\times$ $\checkmark$ $\times$ $\times$
BEHAVIOR [27] I $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
CARLA [14] U $\times$ $\checkmark$ $\times$ 2 $\times$ $\checkmark$ $\checkmark$ $\checkmark$
AirSim [43] U $\times$ $\checkmark$ $\times$ 2 $\times$ $\checkmark$ $\checkmark$ $\checkmark$
MetaUrban [55] U $\checkmark$ $\times$ $\times$ 2 $\checkmark$ $\times$ $\times$ $\checkmark$
EmbodiedCity [18] U $\times$ $\times$ $\times$ 2 $\checkmark$ $\times$ $\times$ $\times$
Grutopia [51] U $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
AerialVLN [30] U $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$